# HEIST: A Graph Foundation Model for Spatial Transcriptomics and Proteomics Data

### v1

@@ 3 Method
HEIST is trained on a combination of contrastive and masked auto-encoding objectives.
Contrastive. We use a contrastive objective to learn context-aware representations by bringing similar cells and genes—such as neighboring cells of the same type or co-expressed genes—closer in embedding space while pushing dissimilar pairs apart. This separates functionally distinct cell populations and gene modules, even when spatially close. Additionally, we introduce cross-level contrastive alignment to ensure consistency between gene and cell representations, so cell embeddings reflect gene expression patterns and gene embeddings incorporate spatial context. The contrastive loss is calculated using the equation below:
$\ell_{c\leftrightarrow c}=\log\frac{e^{\theta({{\mathbf{z}}}_{c,i},{{\mathbf{z}}}_{c,j})/\tau}}{\sum_{k\in{\mathcal{N}}_{i}}e^{\theta({{\mathbf{z}}}_{c,i},{{\mathbf{z}}}_{c,k})/\tau}},\quad\ell_{g\leftrightarrow g}=\log\frac{e^{\theta({{\mathbf{z}}}^{i}_{g,p},{{\mathbf{z}}}^{j}_{g,q})/\tau}}{\sum_{k\in{\mathcal{N}}_{i}}e^{\theta({{\mathbf{z}}}^{i}_{g,p},{{\mathbf{z}}}^{k}_{g,r})/\tau}},\quad\ell_{c\leftrightarrow g}=\log\frac{e^{\theta({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{j}_{g})/\tau}}{\sum_{k\in{\mathcal{N}}_{i}}e^{\theta({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{k}_{g})/\tau}}$
$\mathcal{L}_{\text{contrastive}}=(\sum_{i,j\in\mathcal{P}}\ell_{c\leftrightarrow c}({{\mathbf{z}}}_{c,i},{{\mathbf{z}}}_{c,j})+\ell_{c\leftrightarrow g}({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{j}_{g})+\sum_{(p,q)\in{\mathcal{V}}^{2}}\ell_{g\leftrightarrow g}({{\mathbf{z}}}^{i}_{g,p},{{\mathbf{z}}}^{j}_{g,q}))$
where $\theta(\cdot,\cdot)$ denotes a similarity function (e.g., cosine similarity), $c\leftrightarrow c,g\leftrightarrow g$, and $c\leftrightarrow g$ denotes contrastive loss between cells, genes, and cell-genes respectively, $\tau$ is a temperature parameter that controls the sharpness of the contrastive distribution, and $t_{i}$ is the cell type label of cell $i$. The set of positive pairs $\mathcal{P}=\{(i,j)\mid t_{i}=t_{j},\,d(i,j)\leq r\}$, where $d(i,j)$ is the spatial distance between cells $i$ and $j$, and $r$ is a fixed spatial radius. Similarly, the set of negative samples for cell $i$ $\mathcal{N}_{i}=\{k\mid t_{i}\neq t_{k},\,d(i,k)\leq r\}$, i.e., cells within radius $r$ that belong to a different cell type.
@@
Final loss. Contrastive learning structures the latent space to emphasize biological similarities and differences, promoting better separation of cell types and gene programs. In contrast, masked autoencoding ensures that embeddings retain rich information content necessary for reconstructing gene expression and spatial locations. Together, they prevent trivial or collapsed representations and produce embeddings that are both discriminative and information-rich. Hence, the final objective is a weighted sum of the contrastive and autoencoding losses, along with an orthogonality regularization which encourages the embedding dimensions to be decorrelated, promoting diverse and non-redundant representations [46]:
@@
where $\lambda$ is a regularization weight, and $\gamma$ is a learnable scalar that dynamically balances the contrastive and reconstruction terms.
Positional encoding. HEIST incorporates positional encodings (PE) at both the cell and gene levels to inject spatial and regulatory structure into the learned representations. As opposed to traditional graph PEs such has Laplacian PE [10], random walk PE [10], node-centrality based PE [44], we make use of sinusoidal PEs. This design choice is motivated by the fact that sinusoidal PE yields expressive representations and is computationally inexpensive as opposed to traditional graph PEs, leading to efficient and expressive representations [40]. Specifically, we use 2D-sinusoidal embeddings and rank-based sinusoidal described in the Appendix Section B.1. We then add the spatial information and gene-expression to the positional encodings using the equations below:
where ${\mathbf{H}}_{c}^{(0)}$ and ${\mathbf{H}}_{g}^{k^{(0)}}$ are the input cell and gene embeddings for cell $k$, respectively.
Computational efficiency As shown in Table A4 in the Appendix, HEIST demonstrates significant computational advantages, achieving 8× faster embedding extraction time compared to scGPT-spatial and 48× faster than scFoundation. This efficiency comes from HEIST’s hierarchical modeling and sparse cross-level message passing, which avoids the expensive full self-attention computations required by large language model-based approaches.
@@ 4 Experiments
In this section, we first describe the pretraining datasets used to train HEIST with contrastive and autoencoding objectives. We then outline the downstream tasks and corresponding datasets, followed by baselines, results, insights, and ablation studies.
@@
Pretraining Datasets. HEIST is trained on a large and diverse collection of high-resolution spatial transcriptomics datasets, primarily generated using single-cell technologies such as MERFISH and Xenium. The pretraining dataset comprises 22.3 million cells from 124 tissue slices across 15 organs [cf. Figure 1(A)], including 13.3M cells from 10x Genomics  [^1] https://www.10xgenomics.com/, 8.7M from Vizgen  [^2] https://vizgen.com/, and 360k cells from the Seattle Alzheimer’s Brain Atlas  [^3] https://portal.brain-map.org/explore/seattle-alzheimers-disease. The large scale and diversity of the dataset improve the reliability and robustness of learned representations, enabling better transferability to downstream tasks across varying biological contexts and experimental platforms. While integrating data from different technologies presents challenges such as batch effects and varying gene panels, HEIST’s architecture and training objectives are specifically designed to overcome these issues and learn biologically meaningful representations. A detailed breakdown of datasets, descriptions, and sources is provided in Appendix C.
@@
Experimental setup. We detail pretraining hyperparameters in Appendix Table A1. HEIST was pretrained on 4 NVIDIA L40s GPUs (40GB each), with each epoch taking approximately 3 hours. Although the maximum number of epochs was set to 20, early stopping based on validation loss typically halted training around Epoch 5 or 6. For downstream evaluations, we assess HEIST in both zero-shot and fine-tuning settings. In the zero-shot setting, the pretrained model is directly evaluated on unseen spatial transcriptomics datasets to assess generalization without further training. For fine-tuning, we first extract embeddings from the frozen model and either train an MLP prediction head or fine-tune the decoder. We perform each experiment 5 times provide mean and standard deviation, except in Charville and UPMC dataset where we use the split the data using method from Wu et al. [41]. The code is available at https://anonymous.4open.science/r/HEIST-908D.
@@
Downstream Tasks. We evaluate HEIST across four spatial transcriptomics technologies, six organs, and four downstream tasks—cell clustering, cell type annotation, clinical outcome prediction, and gene imputation—to assess both biological insight discovery and clinical relevance in zero-shot and fine-tuned settings.
@@
Cell clustering is critical for discovering novel cell types and understanding how microenvironmental factors shape cellular behavior, particularly in tumor microenvironments. An expressive model should not only cluster cells but also reveal microenvironment-driven subclusters, providing insights into spatially informed cell populations. Clustering is performed using frozen embeddings, and performance is evaluated on datasets such as SEA-AD [13], Charville, UPMC, DFCI [41], and Merfish Lung Cancer [5] using normalized mutual information (NMI). Cell type annotation classifies cells into known biological categories, enabling functional interpretation of cellular diversity. HEIST embeddings are extracted from the frozen model, and an MLP classifier is trained on labeled datasets including SEA-AD, Charville, UPMC, DFCI, and MERFISH lung cancer, with performance evaluated using F1 score.
@@
Clinical outcome prediction aims to classify entire tissues, predicting outcomes such as immunotherapy response, treatment outcomes, remission status, and placenta condition. This task is essential for clinical decision-making and understanding disease progression. HEIST is evaluated on datasets including Charville (Colon), UPMC (Neck), and DFCI (Neck) [41] collected using CODEX, skin cancer data [30] collected using MIBI, and placenta data collected using Xenium. Predictions are made using frozen cell embeddings and an MLP classifier, with AUC-ROC reported as the evaluation metric. Gene imputation recovers missing or noisy gene expression values, a common issue in spatial transcriptomics due to measurement limitations. We perform gene imputation by predicting masked gene values, using stratified sampling based on gene sparsity following the approach of Avşar and Pir [1]. This task is evaluated in both zero-shot and by fine-tuning the decoder, reporting Pearson correlation between predicted and true gene expression values. We explain these tasks in further detals in Appendix C.2.
Baselines. We compare HEIST with graph-based spatial models STAGATE [9] and GraphST [26], which focus on local spatial relationships but lack generalization across datasets, and with scFoundation [18], a large-scale single-cell foundation model that ignores spatial context. We also include comparisons with recent spatial foundation models CellPLM [39] and scGPT-spatial [38], which incorporate spatial information but do not explicitly model hierarchical gene-cell interactions. These baselines provide a comprehensive evaluation across specialized spatial models, non-spatial foundation models, and spatial foundation models, demonstrating HEIST’s ability to generalize across biological scales and outperform both specialized and general-purpose models.
@@
Figure 3: HEIST accounts for tissue microenvironments. A. Comparison of cell embeddings for the same tissue slice from SEA colored by cell types. HEIST demonstrates superior separation of distinct cell types compared to other methods. B. Spectral clustering of HEIST embeddings reveals sub-clustering within cell types (refer to Appendix  C.2 for more details). C. L4-IT neurons investigation shows that clusters accounts for spatial organization and neighborhood differences, demonstrating that HEIST embeddings effectively captures spatial microenvironments through cross-message passing. Numerical results are available in Table A3 in Appendix Section D.
@@ 4.1 Results
Spatially-aware cell type discovery. Figure 3 shows that HEIST is able to do separate spatially-aware subclusters based on cell-embeddings, while existing foundation models fail to do so. Our model separates known cell types and further reveals microenvironment-driven subpopulations, enabling the discovery of new, spatially informed cell types. Meanwhile, CellPLM an scGPT-spatial fail to address the microenvironment separation as shown in Figure A2 in Appenndix.
Tissue-classification. Table 1 reports the tissue-level classification performance of various models across multiple datasets and tasks, measured by AUC-ROC. HEIST consistently outperforms existing models, achieving the highest scores in six out of seven evaluation scenarios. Notably, on the UPMC tasks, HEIST surpasses the foundation models scGPT-spatial and CellPLM by 25.4% and 30%, respectively. Since CODEX is not a technology that scGPT-spatial and CellPLM were trained on, they support only 6 out of 40 genes in the UPMC datasets. As a result, these models ignore most of the available gene information, leading to significant performance degradation. In contrast, HEIST leverages GRNs and models gene dynamics directly, rather than relying on a fixed set of genes. This enables HEIST to achieve substantially better performance across diverse technologies and datasets.
Organ Colon Neck Neck Skin Placenta
Dataset Charville UPMC DFCI Melanoma
Model Task Outcome Recurrence Outcome Recurrence Outcome Response Condition
STAGATE 0.657 $\pm$ 0.032 0.783 $\pm$ 0.050 0.602 $\pm$ 0.054 0.659 $\pm$ 0.013 0.633 $\pm$ 0.210 0.533 $\pm$ 0.267 0.578 $\pm$ 0.12
Space-gm* 0.793 0.696 0.863 0.883 0.873 - -
[…]
CellPLM 0.744 $\pm$ 0.006 0.801 $\pm$ 0.032 0.681 $\pm$ 0.134 0.667 $\pm$ 0.045 0.750 $\pm$ 0.083 0.580 $\pm$ 0.133 0.682 $\pm$ 0.15
scGPT-spatial 0.834 $\pm$ 0.081 0.806 $\pm$ 0.019 0.717 $\pm$ 0.117 0.676 $\pm$ 0.097 0.875 $\pm$ 0.041 0.600 $\pm$ 0.000 0.602 $\pm$ 0.08
HEIST 0.861 $\pm$ 0.086 0.887 $\pm$ 0.041 0.835 $\pm$ 0.001 0.929 $\pm$ 0.030 0.937 $\pm$ 0.062 0.866 $\pm$ 0.066 0.769 $\pm$ 0.06
HEIST Imp. % 3.2 10.0 -3.3 5.2 7.1 44.3 12.7
Table 1: Performance on clinical outcome prediction. We classify cancer outcome, cancer remission, treatment response and placental conditions. *Results are taken from [41]
Cell-type annotation. Table 2 shows the performance on the cell type annotation task across multiple datasets. HEIST achieves the highest performance in four out of five datasets, with substantial gains in UPMC and DFCI neck datasets (311.3% and 81.9% improvements, respectively). Notably, scGPT-spatial was pretrained on the MERFISH Lung Cancer dataset, explaining its strong performance there, while HEIST remains competitive despite not being trained on that data. These results again highlight the effectiveness and generalizability of HEIST.
Organ Brain Colon Neck Neck Lung
Model SEA-AD Charville UPMC DFCI
STAGATE 0.3304 $\pm$ 0.0625 0.2759 $\pm$ 0.0490 0.0687 $\pm$ 0.0136 0.0685 $\pm$ 0.0213 0.2187 $\pm$ 0.0570
GraphST 0.2296 $\pm$ 0.1772 0.3675 $\pm$ 0.0873 0.0617 $\pm$ 0.0199 0.0577 $\pm$ 0.0261 0.4081 $\pm$ 0.0658
[…]
CellPLM 0.6701 $\pm$ 0.0827 0.4760 $\pm$ 0.0669 0.0563 $\pm$ 0.0212 0.0565 $\pm$ 0.0179 0.5044 $\pm$ 0.1607
scGPT-spatial 0.5907 $\pm$ 0.0029 0.3494 $\pm$ 0.0624 0.0464 $\pm$ 0.0162 0.0618 $\pm$ 0.0163 0.5671 $\pm$ 0.1685
HEIST 0.9953 $\pm$ 0.0158 0.5340 $\pm$ 0.1293 0.2826 $\pm$ 0.0758 0.1124 $\pm$ 0.0521 0.5126 $\pm$ 0.1170
HEIST Imp. % 48.5 12.2 311.3 81.9 -9.6
Table 2: Performance on cell type annotation. Annotations are provided as a feature in each dataset.
@@
Gene imputation. Table 3 reports the performance on the gene imputation task for the placenta and skin datasets. HEIST achieves the best performance after fine-tuning, surpassing all baseline models by 2.5% on the placenta dataset and 9.1% on the skin dataset. Although HEIST’s zero-shot performance is limited by the dataset-specific nature of gene expression patterns, fine-tuning allows the model to effectively adapt by leveraging its hierarchical structure and GRN-based representations, resulting in improved performance.
@@
Ablation of components. Table 4 shows that removing any key component of HEIST leads to performance degradation across tasks, confirming their importance. Hierarchical modeling and spatial information are most critical, with their removal causing the largest drops. Cross-level message passing and contrastive learning significantly improve cell classification, while the masked autoencoder (MAE) is crucial for gene imputation and classification. Orthogonal regularization mainly benefits gene imputation but has less impact on other tasks.

### v2

@@ 2 Method
Contrastive Learning. We use a contrastive objective to learn context-aware representations by bringing similar cells and genes—such as neighboring cells of the same type or co-expressed genes—closer in embedding space while pushing dissimilar pairs apart. This separates functionally distinct cell populations and gene modules, even when spatially close. Additionally, we introduce cross-level contrastive alignment to ensure consistency between gene and cell representations, so cell embeddings reflect gene expression patterns and gene embeddings incorporate spatial context. The contrastive loss is calculated using the equation below:
$\ell_{c\leftrightarrow c}$ $=\log\tfrac{e^{\theta({{\mathbf{z}}}_{c,i},{{\mathbf{z}}}_{c,j})/\tau}}{e^{\theta({{\mathbf{z}}}_{c,i},{{\mathbf{z}}}_{c,j})/\tau}+\sum_{k\in{\mathcal{N}}_{i}}e^{\theta({{\mathbf{z}}}_{c,i},{{\mathbf{z}}}_{c,k})/\tau}},\quad\ell_{g\leftrightarrow g}=\log\tfrac{e^{\theta({{\mathbf{z}}}^{i}_{g,p},{{\mathbf{z}}}^{j}_{g,q})/\tau}}{e^{\theta({{\mathbf{z}}}^{i}_{g,p},{{\mathbf{z}}}^{j}_{g,q})/\tau}+\sum_{k\in{\mathcal{N}}_{i}}e^{\theta({{\mathbf{z}}}^{i}_{g,p},{{\mathbf{z}}}^{k}_{g,r})/\tau}},$
$\ell_{c\leftrightarrow g}$ $=\log\tfrac{e^{\theta({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{j}_{g})/\tau}}{e^{\theta({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{j}_{g})/\tau}+\sum_{k\in{\mathcal{N}}_{i}}e^{\theta({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{k}_{g})/\tau}},$
$\mathcal{L}_{\text{contrastive}}$ $=\sum_{i,j\in{\mathcal{P}}}\!\big(\ell_{c\leftrightarrow c}({{\mathbf{z}}}_{c,i},{{\mathbf{z}}}_{c,j})+\ell_{c\leftrightarrow g}({{\mathbf{z}}}_{c,i},\bar{{{\mathbf{z}}}}^{j}_{g})\big)+\sum_{(p,q)\in{\mathcal{V}}^{2}}\!\ell_{g\leftrightarrow g}({{\mathbf{z}}}^{i}_{g,p},{{\mathbf{z}}}^{j}_{g,q}).$
where $\theta(\cdot,\cdot)$ denotes a similarity function (e.g., cosine similarity), $c\leftrightarrow c,g\leftrightarrow g$, and $c\leftrightarrow g$ denotes contrastive loss between cells, genes, and cell-genes respectively, $\tau$ is a temperature parameter that controls the sharpness of the contrastive distribution, and $t_{i}$ is the cell type label of cell $i$. The set of positive pairs $\mathcal{P}=\{(i,j)\mid t_{i}=t_{j},\,d(i,j)\leq r\}$, where $d(i,j)$ is the spatial distance between cells $i$ and $j$, and $r$ is a fixed spatial radius. Similarly, the set of negative samples for cell $i$, $\mathcal{N}_{i}=\{k\mid t_{i}\neq t_{k},\,d(i,k)\leq r\}$, i.e., cells within radius $r$ that belong to a different cell type.
@@
Final loss. Contrastive learning structures the latent space to emphasize biological similarities and differences, promoting better separation of cell types and gene programs. In contrast, masked autoencoding ensures that embeddings retain rich information content necessary for reconstructing gene expression and spatial locations. Together, they prevent trivial or collapsed representations and produce embeddings that are both discriminative and information-rich. Hence, the final objective is a weighted sum of the contrastive and autoencoding losses, along with an orthogonality regularization which encourages the embedding dimensions to be decorrelated, promoting diverse and non-redundant representations (Zhang et al., 2021):
@@
where $\lambda$ is a regularization weight, and $\gamma$ is a learnable scalar that dynamically balances two terms.
Computational efficiency. As shown in Table A4 in the Appendix, HEIST demonstrates significant computational advantages, achieving 8× faster embedding extraction time compared to scGPT-spatial and 48× faster than scFoundation. This efficiency comes from HEIST’s sparse modeling, which avoids the expensive full self-attention computations required by transformer based models like scGPT-spatial.
@@ 3 Experiments
In this section, we first describe the pretraining datasets used to train HEIST. We then outline the downstream tasks and corresponding datasets, followed by baselines, results, insights, and ablation studies.
@@
Pretraining Datasets. HEIST is trained on a large and diverse collection of high-resolution spatial transcriptomics datasets, primarily generated using single-cell technologies such as MERFISH and Xenium. The pretraining dataset comprises 22.3M cells from 124 tissue slices across 15 organs [cf. Figure 1(A)], including 13.3M cells from 10x Genomics  [^1] https://www.10xgenomics.com/, 8.7M from Vizgen  [^2] https://vizgen.com/, and 360K cells from the Seattle Alzheimer’s Brain Atlas  [^3] https://portal.brain-map.org/explore/seattle-alzheimers-disease. The large scale and diversity of the dataset improve the reliability and robustness of learned representations, enabling better transferability to downstream tasks across varying biological contexts and technologies. A detailed breakdown of datasets, descriptions, and sources is provided in Appendix D, and we provide a spreadsheet of the dataset details in the supplementary material.
@@
Experimental setup. We provide pretraining hyperparameters in Appendix Table A1. HEIST was pretrained on 4 NVIDIA L40s GPUs (40GB each), with each epoch taking approximately 3 hours. Although the maximum number of epochs was set to 20, early stopping based on validation loss typically halted training around Epoch 5 or 6. For downstream evaluations, we assess HEIST in both zero-shot and fine-tuning settings. In the zero-shot setting, the pretrained model is directly evaluated on unseen spatial transcriptomics datasets to assess generalization without further training. For fine-tuning, we first extract embeddings from the frozen model and either train an MLP prediction head or fine-tune the decoder. We perform each experiment 5 times provide mean and standard deviation, except in Charville and UPMC datasets where we use the split the data using method from (Wu et al., 2022). The code is available at https://anonymous.4open.science/r/HEIST-E3F8.
@@
Downstream Tasks. We evaluate HEIST across four different spatial transcriptomics and proteomics technologies, five organs, and four downstream tasks—cell clustering, cell type annotation, clinical outcome prediction, and gene imputation—to assess both biological insight discovery and clinical relevance of HEIST.
@@
Cell clustering is critical for discovering novel cell types and understanding how microenvironmental factors shape cellular behavior, particularly in tumor microenvironments. An expressive model should not only cluster cells but also reveal microenvironment-driven subclusters, providing insights into spatially informed cell populations. Clustering is performed using frozen embeddings, and performance is evaluated on datasets two spatial transcriptomics datasets SEA-AD (Gabitto et al., 2024) and Merfish Lung Cancer (Chen et al., 2024), and three proteomics datasets Charville, UPMC, and DFCI (Wu et al., 2022), using normalized mutual information (NMI). Cell type annotation classifies cells into known biological categories, enabling functional interpretation of cellular diversity. HEIST embeddings are extracted from the frozen encoder, and an MLP classifier is trained on labeled datasets including SEA-AD, Charville, UPMC, DFCI, and MERFISH lung cancer, with performance evaluated using F1 score.
@@
Clinical outcome prediction aims to classify entire tissues, predicting outcomes such as immunotherapy response, treatment outcomes, remission status, and placenta condition. This task is essential for clinical decision-making and understanding disease progression. HEIST is evaluated on datasets including proteomics data Charville (Colon), UPMC (Neck), and DFCI (Neck) (Wu et al., 2022) collected using CODEX, skin cancer data (Ptacek et al., 2021) collected using MIBI, and spatial transcriptomics data of placenta collected using Xenium. Predictions are made using frozen cell embeddings and an MLP classifier, and we report AUC-ROC. Gene imputation recovers missing or noisy gene expression values, a common issue in spatial transcriptomics due to measurement limitations. We perform gene imputation by predicting masked gene values, using stratified sampling based on gene sparsity following the approach of Avşar & Pir (2023). This task is evaluated in both zero-shot and by fine-tuning the decoder, reporting Pearson correlation between predicted and true gene expression values. We explain these tasks in further details in Appendix D.2.
Baselines. We benchmark HEIST against a diverse set of baselines, including graph-based spatial models, single-cell foundation models, and recent spatial foundation approaches. In particular, we include STAGATE (Dong & Zhang, 2022) and GraphST (Long et al., 2023), which capture local spatial relationships but struggle to generalize across datasets. We also evaluate against scFoundation (Hao et al., 2023), a large-scale single-cell foundation model that ignores spatial context. We also include comparisons with recent spatial foundation models CellPLM (Wen et al., 2023a), NoVAE (Blampey et al., 2024) and scGPT-spatial (Wang et al., 2025), which incorporate spatial information but do not explicitly model hierarchical gene-cell interactions. Finally, for the gene imputation task, we compared against MAGIC a denoising method based on graph diffusion. For proteomics datasets such as UPMC, baselines like scGPT-spatial and CellPLM were originally pretrained on fixed gene vocabularies, supporting only 6 of the 28 markers available. To enable comparison, we report two settings: (i) unaligned, where models process only the subset of overlapping genes, and (ii) aligned, where we manually remap marker identities to the closest supported genes following prior work. It is important to note that scGPT pretrains on single-cell spatial trasncritomics data mixed with Visium (spot-level), while CellPLM combines scRNA-seq with single-cell spatial transcriptomics. HEIST, in contrast, is trained exclusively on single-cell resolution spatial transcriptomics. However, there is major overlap in the single-cell spatial transcriptomics portion of training data between HEIST and the baselines.
@@
Figure 3: HEIST accounts for tissue microenvironments. (A) Comparison of cell embeddings (PHATE) for the same tissue slice from SEA colored by annotated cell types. HEIST demonstrates superior separation compared to other methods. (B) Spectral clustering reveals three spatially-informed sub clusters in L4-IT cells. (C) Visualizing the sub clusters neighborhood distribution shows that each cluster accounts for neighborhood differences, demonstrating that HEIST embeddings effectively captures spatial microenvironments through cross-message passing. We show similar results for CEllPLM and scGPT-spatial in Figure A2, showing that these models can not infer spatially-informed clusters. Numerical results are available in Table A3 in Appendix Section E. (D) Silhouette Score comparision between each method.
@@ 3.1 Results
Spatially-aware cell type discovery. A key strength of HEIST is its ability to uncover spatially-informed cellular subpopulations. Figure 3 shows that HEIST embeddings not only separate canonical cell types, but also resolve finer subclusters that align with local microenvironmental context. These subclusters correspond to meaningful biological distinctions. Existing approaches like CellPLM and scGPT-spatial fail to capture sub clusters and collapse such substructure, missing microenvironment effects (Figure A2, Appendix). This ability to differentiate spatially-driven heterogeneity is essential for discovering novel biomarkers.
Table 1: Performance on gene imputation task
@@
Gene imputation. Table 1 reports the performance on the gene imputation task for the placenta and skin datasets. HEIST achieves the best performance after fine-tuning, surpassing all baseline models by 2.5% on the placenta dataset and 9% on the skin dataset. We highlight comparisons with MAGIC, as our preprocessing also incorporates this method. The improvements achieved by HEIST can be attributed to its cross-attention mechanism, which integrates contextual information from neighboring cells to enhance imputation for each target cell. Although HEIST’s zero-shot performance is limited by the dataset-specific nature of gene expression patterns, fine-tuning allows the model to adapt effectively by leveraging its hierarchical structure and gene co-expression networks, resulting in improved performance.
Table 2: Performance on clinical outcome prediction. We classify cancer outcome, cancer remission, treatment response and placental conditions. Results on the right of vertical are proteomics. NEM stands for "not enough markers" error from Novae when it can not match enough markers. *Results from (Wu et al., 2022)
Organ Placenta Colon Neck Neck Skin
Dataset Charville UPMC DFCI Melanoma
Model Task Condition Outcome Recurrence Outcome Recurrence Outcome Response
STAGATE 0.578 $\pm$ 0.12 0.657 $\pm$ 0.032 0.783 $\pm$ 0.050 0.602 $\pm$ 0.054 0.659 $\pm$ 0.013 0.633 $\pm$ 0.210 0.533 $\pm$ 0.267
Space-gm* - 0.793 0.696 0.863 0.883 0.873 -
ScFoundation 0.601 $\pm$ 0.16 0.713 $\pm$ 0.122 0.787 $\pm$ 0.113 0.678 $\pm$ 0.061 0.689 $\pm$ 0.111 0.742 $\pm$ 0.093 0.500 $\pm$ 0.155
Novae 0.619 $\pm$ 0.10 0.739 $\pm$ 0.006 0.500 $\pm$ 0.000 NEM NEM 0.750 $\pm$ 0.095 NEM
CellPLM (unaligned) 0.682 $\pm$ 0.15 0.744 $\pm$ 0.006 0.801 $\pm$ 0.032 0.681 $\pm$ 0.134 0.667 $\pm$ 0.045 0.750 $\pm$ 0.083 0.580 $\pm$ 0.133
CellPLM (aligned) - 0.732 $\pm$ 0.007 0.792 $\pm$ 0.030 0.670 $\pm$ 0.128 0.655 $\pm$ 0.043 0.738 $\pm$ 0.080 -
scGPT-spatial (unaligned) 0.602 $\pm$ 0.08 0.834 $\pm$ 0.081 0.806 $\pm$ 0.019 0.717 $\pm$ 0.117 0.676 $\pm$ 0.097 0.875 $\pm$ 0.040 0.600 $\pm$ 0.000
scGPT-spatial (aligned) 0.594 $\pm$ 0.081 0.854 $\pm$ 0.005 0.702 $\pm$ 0.001 0.789 $\pm$ 0.080 0.858 $\pm$ 0.006
HEIST 0.769 $\pm$ 0.06 0.861 $\pm$ 0.086 0.887 $\pm$ 0.041 0.835 $\pm$ 0.001 0.929 $\pm$ 0.030 0.937 $\pm$ 0.062 0.866 $\pm$ 0.066
HEIST Imp.% 12.7 3.2 3.5 -3.3 5.2 7.1 44.3
Tissue-classification. Table 2 reports the tissue-level classification performance of various models across multiple datasets and tasks, measured by AUC-ROC. HEIST consistently outperforms existing models, achieving the highest scores in six out of seven evaluation scenarios. Notably, on the UPMC tasks, HEIST surpasses the foundation models scGPT-spatial and CellPLM by 25.4% and 30%, respectively due to scGPT-spatial and CellPLM ignoring most of the available markers. Even after aligning, performance remains limited because these models cannot fully leverage all the available markers. In contrast, HEIST constructs co-expression networks directly from observed proteins, allowing it to incorporate all available markers without retraining. This structural flexibility underlies its strong generalization to proteomics.
Cell-type annotation. Table 3 shows the performance on the cell type annotation task across multiple datasets. HEIST achieves the highest performance in four out of five datasets, with substantial gains in UPMC and DFCI neck datasets (28.7% and 17.9% improvements, respectively). Notably, scGPT-spatial was pretrained on the MERFISH Lung Cancer dataset, explaining its strong performance there on that data. These results again highlight the effectiveness and generalizability of HEIST.
Table 3: Performance on cell type annotation. Annotations are provided as a feature in each dataset. Results on the right of vertical are proteomics.
Organ Lung Brain Colon Neck Neck
Model SEA-AD Charville UPMC DFCI
STAGATE 0.2187 $\pm$ 0.0570 0.3304 $\pm$ 0.0625 0.2759 $\pm$ 0.0490 0.0687 $\pm$ 0.0136 0.0685 $\pm$ 0.0213
GraphST 0.4081 $\pm$ 0.0658 0.2296 $\pm$ 0.1772 0.3675 $\pm$ 0.0873 0.0617 $\pm$ 0.0199 0.0577 $\pm$ 0.0261
ScFoundation 0.150 $\pm$ 0.014 0.2495 $\pm$ 0.1147 0.3220 $\pm$ 0.1421 0.0222 $\pm$ 0.0079 0.041 $\pm$ 0.021
Novae NEM 0.2332 $\pm$ 0.0434 0.2194 $\pm$ 0.0455 NEM 0.0736 $\pm$ 0.0215
CellPLM (unaligned) 0.5044 $\pm$ 0.1607 0.6701 $\pm$ 0.0827 0.4760 $\pm$ 0.0669 0.0563 $\pm$ 0.0212 0.0565 $\pm$ 0.0179
CellPLM (aligned) - - 0.3047 $\pm$ 0.0040 0.0337 $\pm$ 0.0032 0.0413 $\pm$ 0.0015
scGPT-spatial (unaligned) 0.5671 $\pm$ 0.1685 0.5907 $\pm$ 0.0029 0.3494 $\pm$ 0.0624 0.0464 $\pm$ 0.0162 0.0618 $\pm$ 0.0163
scGPT-spatial (aligned) 0.3280 $\pm$ 0.0499 0.2195 $\pm$ 0.0490 0.0953 ± 0.0190
HEIST 0.5126 $\pm$ 0.1170 0.9953 $\pm$ 0.0158 0.5340 $\pm$ 0.1293 0.2826 $\pm$ 0.0758 0.1124 $\pm$ 0.0521
HEIST Imp. -9.6 % 48.5 12.2 28.7 17.9
@@
Ablation of components. Table 4 shows that removing any key component of HEIST leads to performance degradation across tasks, confirming their importance. Hierarchical modeling and spatial information are most critical, with their removal causing the largest drops. Pre-training is crucial, especially for clinical outcome prediction due to the skewed label distribution in these datasets. Cross-level message passing and contrastive learning significantly improve cell classification, while the masked autoencoder (MAE) is crucial for gene imputation and classification.